Circular No. 7662

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dc IN NGC 5777 G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong (Rolvenden, Kent) of an apparent supernova at α = 14^{h}51^{m}16^{s}.15, δ = +58^{o}59′02′′.8 (equinox 2000.0; average from three CCD images obtained with a 0.35-m reflector), said to be 26′′ west and 28′′ north of the center of NGC 5777. Reported magnitudes for SN 2001dc [unfiltered CCD data unless otherwise noted, and from Armstrong unless otherwise noted; ‘–’ means not visible; JKT = Jacobus Kapteyn Telescope (via P. Meikle, Imperial College; uncertainty ± 0.1 mag; June 22 data obtained by D. Bramich, R. Corradi, and P. Erwin)]: 1995 Apr. 4 UT, [20.8 (Palomar Sky Survey red plate); Apr. 18, [21 (JKT I-band images); Apr. 24, [22.5 (PSS blue plate); 1997 Apr. 18, – (PSS infrared plate); 2000 Dec. 3, –; 2001 Apr. 12, –; May 12, –; 21.977, 18.6 (prediscovery); 30.958, 18.5 (discovery; limiting mag 19.3); June 1.021, 18.5; 2.985, 18.3 (T. Boles, Coddenham, England); 22.93, B = 21.11, V = 19.83, R = 19.15, I = 18.46 (JKT). M. Irwin (Cambridge Astronomical Survey Unit) adds that a June 22.93 JKT image yields position end figures 16^{s}.21, 03′′.0 for SN 2001dc, which is ∼ 14′′ west and 22′′ north of the nucleus of NGC 5777. Meikle notes that JKT data show that the change in brightness of SN 2001dc from June 22.93 to July 4.92 was < 0.3 mag in B, 0.1 mag in V, and 0.05 mag in R (all 1σ confidence); there was marginal evidence of a brightening in I (by −0.12 ± 0.04). Meikle and A. Fassia add: “A spectrum obtained on July 10.97 at the Isaac Newton Telescope shows a strong but quite narrow P-Cyg profile in Hα. Together with the lack of significant decline in optical brightness between June 22 and July 4, we conclude that this is a type-IIp supernova, still in its plateau phase during this period. Other lines identified include Fe II (501.8 and 516.9 nm), Na I D, and the Ca II infrared triplet. A blueshifted absorption feature due to Hβ may also be present, but it is close to the blue limit of the spectrum (482 nm). The trough of the Hα profile is blueshifted by ∼ 2400 km/s with respect to the emission peak. The corresponding figure for Na I D is 2300 km/s. The supernova appears to be an unusually low-energy event. Assuming negligible extinction, the V -band magnitude corresponds to M_{V} = −12.6 (adopting 32.4 as the distance modulus for NGC 5777), compared with more typical values (M_{V} = −15 to −16) for a type-IIp event during its plateau phase (Patat et al. 1994, A.Ap. 282, 731). Moreover, the velocity seen in Hα is only about half of that normally observed during this era. Further photometric and spectroscopic observations are strongly encouraged. Of particular interest is the size of the decline immediately following the end of the plateau phase.”

2001 July 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green